Project summaries. ALICE Management Committee meeting : 27 July (8 August) 2012.

frontdotardUrban and Civil

Nov 15, 2013 (3 years and 4 months ago)

98 views

Project summaries.

ALICE Management Committee meeting :
27 July

(8 August)
2012.


Machine development

Photoelectron Gun was successfully HV conditioned to >
400kV after cathode change and
gun bakeout. Bare wafer had no field emission at the test voltage of 370kV. The first cathode
activation achieved just under 3% QE.

Cathode lifetime is reasonable : re
-
Caesiation is normally required once a week. But the gun

vacuum is low 10
-
11
mbar (an order of magnitude higher than we used to have) due to a failure
of one of the NEGs.

First use of INJ quads degausser had a huge effect on beam optics in ALICE injector thus
necessitating re
-
tuning all setups.
Work on various
setups continued.

New AP setup at 60pC have been largely established that could be used for both transverse
and longitudinal beam dynamics experiments.

Towards the end of July, FEL stopped lasing

for currently (as of 27/07/2012) unknown
reasons;
investig
ation continues.

Several machine related issues have been investigated: stability (while there was a period of
large beam instability in July); beam loading effects.

New EMMA style electronics was tested on ALICE and proved to be a very useful tool.


AP
: transverse dynamics

Transverse dynamics studies on ALICE have concentrated on further analysis work using
quadrupole scans around the machine. Work was performed in both the injector and post
-
linac sections of the machine to characterise the transverse d
ynamics, and to attempt to
understand the results previously seen.

The work is ongoing, but has demonstrated the ability to characterise and correct transverse
optics in the post
-
linac section. This is especially important in the AEMITER working mode,
wher
e the low bunch charge of ~20pC enables a more Guassian beam shape, improving the
analysis, but which has much stricter requirements on beam optics in the interaction and
detection regions.

Transverse dynamics in the injector area have also been analysed i
n terms of emittance, and
show the expected behaviour versus buncher power. These studies have demonstrated the
benefits of the new 325keV working point.

A major component of the ongoing work is to understand the effects of bunch charge on
transverse dynam
ics, and this has been investigated thoroughly in terms of the 2
-
beam
dynamics studies. This novel analysis has investigated the high bunch charge, strong
bunching, effects that lead to a double
-
beam, with different transverse properties, throughout
the ma
chine.


Transverse optics set
-
ups for the various working modes of ALICE, have been successful
throughout the current period. Standard set
-
ups for AEMITER, FELIS, AP and THz modes
have all been developed, and show relatively consistent behaviour over the p
eriod. Limiting
factors for the transverse dynamics repeatability are related to the remnant fields in the
magnets, as well as large variability in the longitudinal dynamics which necessitate new
transverse dynamics modifications.



AP:
longitudinal

dynami
cs

Bunch evolution in injector. Some linac
-
exit bunch length measurements were taken in the
‘set
-
up’ period in the final week of June, using RF streak methods. This confirmed the
observations earlier in the year that the booster chirp seemed to have only s
mall effect on the
linac
-
exit bunch length for the current injector set
-
up. This is somewhat unexpected and
further data/simulations are required to understand the effect.



Path length Measurements. Time of flight measurements have continued using high
r
esolution oscilloscope and BPM signals. These have demonstrated the ability to measure
TOF differences to a couple of pico seconds and thus can measure in principle

R56 values of
~ cm level. The effect of lattice tuning including arc quadrupole and sextup
le strength on the
time of flight has been measured, but the repeatability of the results and comparison with
theoretical expectation is not clear and work is underway to understand the experimental
artefacts involved in the method. Elegant simulations hav
e shows some agreement

with the
data on the shape of the TOF vs energy curve and the influence of arc sextupole.



Single Bunch BPMs. Primarily to aid the understanding of the path length measurements,
single bunch BPM electronics have been used on the ar
c BPMs. These have provided
information on the energy of the bunches over the train and the alignment of the magnets in
the arc. In particular they are useful to understand the offset of the beam in the sextupoles
which is crucial to longitudinal dynamics.




2 scheduled AP shifts were lost to RF faults on Sun 1
st

of July.


FEL

FEL lasing was restored very early in the running period after the cathode change (in the first
full shift). A reasonable setup was established with similar power levels to previous
running
periods, however the performance in terms of stability has been mixed


a few shifts with
high stability but more often not so good. This has been a cross
-
machine problem and the
FEL has been used to investigate the source of the instability, in ad
dition to working to
improve the tolerance of the FEL setup to instability. Further work is required on improving
the setup


particularly since some recent changes have made restoring lasing problematic.


A fast
-
response detector was installed in the acce
lerator hall and has been used to resolve
individual pulses in the FEL output. Data was recorded to evaluate the gain of the FEL, with
the intention for this detector to be used in future as an online diagnostic of FEL
performance. A laser used for monitor
ing the FEL cavity mirror angles had failed and has
been replaced.


A new aim for the project is to demonstrate lasing at 20um (a particularly useful wavelength
for scientific studies using the output), which requires a


new electron beam setup at lower
en
ergy. A few attempts have been scheduled but not completed due to various difficulties, so
this remains a priority for future shifts. A detector sensitive to this wavelength has been
sourced and installed.


One other significant development is that several

non
-
FEL
-
specialist members of the
commissioning team have successfully established lasing in this period. This increases the
potential for utilising the FEL in conjunction with other accelerator physics projects, plus
providing additional support for FEL
-
user projects such as FELIS.


EMMA

"EMMA had beam time for three days from 16 to 18 July. A beam in EMMA seemed very
good or even better than ever before probably because of careful tuning by JJ and/or more
stable operation of ALICE. We spent shifts for tw
o subjects. One is the investigation of
EPICS/BPM system and the other is the COD analysis. Unfortunately, the EPICS/BPM
system is still not fully functioning. In order to pin down where the faults occurs, GC and
Richard Burrell who is a designer of the di
gital system from a company spent large fraction
of time. It suggested some improvement scenario to be tested later. Regarding the COD
analysis, DK, IK and JG took a response function by moving

a couple of magnets
transversely. Hopefully, the detailed ana
lysis of this pilot measurement will fix the
parameters to be used for the response function measurement of the entire magnets."


AEMITR


Last year, we reached the stage of running some full system tests (including THz generation
in the accelerator area, a
chieving the required electron beam optics, and operating all the
diagnostics).


The results we obtained showed (as we suspected) a level of beam size jitter
that could make it difficult to identify the signal we are looking for (a variation in beam size
i
n the chicane, indicating an increased energy spread resulting from interaction with the THz
pulse).


Although we could in principle average over many pulses to try to improve the
effective signal
-
to
-
noise ratio, the problem is that our previous timing tec
hnique only enabled
us to synchronise the electron and THz pulses to within (roughly) a nanosecond, whereas the
electron beam is a couple of picoseconds, and the THz pulse perhaps less than that.


This
means we need to make long timing scans while looking
for a signal, which already takes a
considerable time, even without averaging.


To try to improve our sensitivity to what we expect to be a rather small signal, we are
working on two significant modifications to the set
-
up:


1.


Using fast gated cameras in

the chicane, we should be able to record YAG images from
two bunches within the same train.


We expect that the variation in beam size between
bunches in the same train should be significantly smaller than the variation in bunches from
different trains; t
herefore, by looking for differences between bunches within the same train
we should achieve much better sensitivity to the THz interaction.


2.


Use of the streak camera to synchronise the THz and electron bunches to within a few
picoseconds.


This will e
nable us to concentrate our search in regions where we have a much
better confidence that the interaction should be.


In the shifts we have had so far this year, the operation of the gated cameras for observing
different bunches within a single train has b
een demonstrated.


Initial attempts to use the
streak camera identified some optics issues (to do with directing and focusing the image from
the OTR screen onto the camera), which should be resolved with some further work.


At the
same time, continued expe
riments with THz production in the Diagnostics Room have given
a better understanding of the properties of the THz beam, including the focusing required to
produce an observable effect on the electron beam.


The goal is to demonstrate an energy modulation
on the electron beam by interaction with the
THz pulse, by the end of this year.


FELIS

First snom beamtime in July allowed re
-
installing the microscope and testing improved data
acquisition systems which include incident flux normalisation. Some beam inst
ability in the
first part of the beamtime gave opportunity for this to be tested and the expected
improvement in snom images were obtained by normalising to the incident flux. A short
amount of time was used to try to improve the coupling of light from the

fibre onto the MCT
detector element but this will now be perused off line using a broadband thermal source. The
SNOM tip was used to collect the reflected IR light as a function of wavelength for standard
samples and for one tissue sample at various poin
ts on each of the samples. The local
absorption is a small component of the total snom signal and a better understanding of basic
physics of the snom process will now be a major component of a more focussed EPSRC
proposal by PW.


THz

High levels of THz

were generated and transmitted to the diagnostics room during start up.
Although beamline transmission dropped for the exploitation slot on 15 July, Martin Volk
was still able to obtain data from peptides in solution as a function of temperature, using a

custom cell with TPX windows. Another microfluidic cell equipped with silicon windows
was tested but found to cut out much more light. This will now be modify and the
experiment re
-
run on 28 July. A coupler on the THz M1 mirror drive has failed, preve
nting
use of the AP THz detector. This has now been repaired and the beamline will be re
-
optimised on 28 July prior to THz experiments.


THz CTR

A small nickel coated silicon screen is in place in the machine and a small breadboard for test
detectors in place for initial tests of THz generation by CTR. A larger screen and appropriate
holder are now available for future installation.